# On the Effectiveness of ASR Representations in Real-world Noisy Speech Emotion Recognition

### v3

@@ V-B Effect of Different ASR Representations
A meticulous comparative analysis was conducted to assess the impact of various ASR representation approaches on the task of NSER, focusing on different layers within the ASR encoder and decoder. The results of this examination are systematically shown in Table II, which illustrates the performance variations across distinct ASR representation strategies (Last, Mean, and Adapter) within both the encoder and decoder.
The results provide evidence for the effectiveness of our proposed method (Encoder + Decoder), which integrates representations from both the encoder and decoder through the adapter method. This approach consistently achieves better UAR and F1 performances across all datasets and noise conditions, surpassing standalone encoder and decoder methods.
For the MELD dataset, compared with using only the Encoder with the Adapter, the proposed approach yields improvements of 2.24 in UAR and 0.86 in F1. Similarly, compared with using only the Decoder with the Adapter, it provides gains of 2.58 in UAR and 1.32 in F1. These results highlight the importance of jointly leveraging representations from both components to maximize performance.
On the IEMOCAP dataset, the proposed approach surpasses the standalone Encoder with the Adapter by 1.48 in UAR and 2.43 in F1 and the standalone Decoder with the Adapter by 3.75 in UAR and 4.14 in F1. The improvements underscore the advantage of fusing encoder and decoder representations in clean recording environments.
For the IEMOCAP-Environment subset under noisy conditions, the proposed approach yields increases of 2.44 in UAR and 2.99 in F1 compared with the Encoder alone with the Adapter, and 3.28 in UAR and 4.03 in F1 compared with the Decoder alone with the Adapter. The capability to handle environmental noise more effectively demonstrates the robustness of our proposed design.
In the IEMOCAP-Human subset, which contains the most challenging noise conditions, the proposed approach shows improvements of 1.75 in UAR and 2.03 in F1 over the Encoder alone with the Adapter, and 1.80 in UAR and 1.78 in F1 over the Decoder alone with the Adapter. These results indicate that integrating encoder and decoder representations enables the proposed method to better address the complexities of human-induced noise.
Overall, our proposed approach consistently demonstrates superior performance by integrating features from both the encoder and decoder via the Adapter method. This strategy captures complementary information across multiple layers of the ASR architecture, leading to significant gains in NSER performance under a variety of datasets and noise conditions. The findings highlight the critical importance of leveraging both representations for emotion recognition tasks.
V-C Robustness of ASR Representations on NSER
To evaluate the robustness of ASR representations for noisy emotion recognition, we conducted a comparative analysis of various ASR representations, including the encoder with the Adapter (Whisper-Encoder), the decoder with the Adapter (Whisper-Decoder), and a combination of both encoder and decoder with the Adapter (Whisper-Encoder+Decoder), alongside SSL model representations, across various noise intensities. The results are summarized in Table III.
| Model |
| SNR -5 | SNR 0 | SNR 5 | SNR -5 | SNR 0 | SNR 5 | |||||||||||
| UAR (%) | F1 (%) | UAR (%) | F1 (%) | UAR (%) | F1 (%) | UAR (%) | F1 (%) | UAR (%) | F1 (%) | UAR (%) | F1 (%) | |||||
| Wav2Vec 2.0 | 57.35 | 56.41 | 61.92 | 60.32 | 65.02 | 64.19 | 50.26 | 49.26 | 59.77 | 57.85 | 64.40 | 62.66 | ||||
| Hubert | 57.71 | 56.21 | 61.24 | 60.26 | 64.53 | 63.94 | 54.62 | 54.57 | 60.49 | 60.02 | 64.40 | 62.90 | ||||
| WavLM | 56.17 | 54.40 | 60.31 | 58.49 | 63.24 | 62.49 | 56.63 | 54.33 | 59.10 | 57.01 | 61.65 | 60.89 | ||||
| Whisper-Encoder | 68.36 | 67.58 | 70.56 | 69.43 | 73.14 | 72.66 | 61.45 | 60.77 | 67.04 | 66.36 | 70.08 | 70.22 | ||||
| Whisper-Decoder | 67.20 | 65.99 | 69.72 | 68.39 | 70.07 | 68.99 | 60.42 | 60.19 | 66.99 | 66.61 | 68.38 | 68.78 | ||||
| Whisper-Encoder+Decoder | 68.78 | 66.34 | 73.00 | 72.42 | 73.84 | 73.41 | 63.06 | 61.58 | 68.79 | 68.39 | 70.81 | 70.71 | ||||
The results reveal that as noise intensity increases, the performance of all models declines, with human noise causing greater fluctuations than environmental noise. Conversely, reducing noise intensity leads to consistent performance improvements across the board, demonstrating the robustness of ASR representations.
In the Environment subset of IEMOCAP, the performance of the Whisper-Encoder+Decoder decreases as noise intensifies. When the SNR drops from 0 to -5 dB, the UAR decreases by 4.22, whereas the F1 score declines by 6.08. Conversely, an increase in SNR from 0 to 5 dB results in improvements of 0.84 in UAR and 0.99 in F1. These results demonstrate that although performance is affected under severe noise conditions, the Whisper-Encoder+Decoder maintains notable robustness. Compared with the best-performing SSL representation (Wav2vec 2.0), the Whisper-Encoder+Decoder achieves significant improvements, with gains of up to 11.08 in UAR and 12.11 in F1 at 0 dB. Furthermore, although the Whisper-Encoder shows slightly better stability than the Whisper-Decoder under challenging noise conditions, the integration of both representations ensures the most consistent and robust performance across all SNR levels.
In the Human subset, which presents more complex human noise challenges, performance declines sharply with increasing noise intensity. For the Whisper-Encoder+Decoder, the UAR decreases by 5.73 and the F1 score drops by 6.81 as the SNR decreases from 0 to -5 dB. Conversely, when the SNR increases to 5 dB, the Whisper-Encoder+Decoder achieves improvements of 2.02 in UAR and 2.32 in F1 relative to 0 dB. Compared with the best-performing SSL representation (Hubert), the Whisper-Encoder+Decoder achieves significant improvements, with gains of up to 8.3 in UAR and 8.37 in F1 at 0 dB. This highlights the superior capability of the Whisper-Encoder+Decoder to address complex human noise effectively.
Overall, the findings confirm the superior robustness of ASR representations, particularly under noisy conditions. Compared with SSL-based models, the proposed method (Whisper-Encoder+Decoder) achieves consistently better performance across all noise levels, with significant improvements observed in both environmental and human noise scenarios.
V-D Correlation between ASR Performance and NSER Performance
@@
For the MELD dataset, tested under mixed real-world noise conditions, Whisper achieves a WER of 0.52 and an F1 of 48.87 for NSER. Conversely, Conformer records a WER of 0.75 and an NSER of 40.27 for F1. For the IEMOCAP dataset, Whisper attains a WER of 0.28 and an NSER of 76.53 for F1. On the other hand, Conformer exhibits slightly higher values, with a WER of 0.55 and an NSER of 73.04 for F1.
In the IEMOCAP-Environment subset, the ASR performance gradually deteriorates as the SNR decreases. For Whisper, at an SNR of 5 dB, the WER increases to 0.32. At an SNR of 0 dB, this metric rises to 0.39. Under the most challenging condition of -5 dB, the WER peaks at 0.54. Additionally, the F1 score for NSER tasks decreases by 7.07 as the noise level increases from 5 to -5 dB, indicating degraded performance in NSER tasks under higher noise conditions. For Conformer, performance shows even greater sensitivity to noise, with WER increasing to 0.65 at 5 dB SNR, rising further to 0.75 at 0 dB SNR, and reaching 0.86 at -5 dB SNR. The F1 score for NSER using the Conformer model decreases by 6.01 as the noise level increases from 5 to -5 dB, indicating that the Conformer model’s performance deteriorates under higher-noise-level conditions.
@@
In the IEMOCAP-Human subset, performance degradation is more pronounced. For the Whisper model, at SNR of 5 dB, the WER is 0.36. This WER increases to 0.57 at SNR of 0 dB and further increases to 0.86 at -5 dB, indicating a substantial decline in performance with a decreasing SNR. Concurrently, the F1 score for NSER decreases by 9.13 as the noise level increases from 5 to -5 dB. For the Conformer model, performance worsens consistently across all noise levels, with the WER increasing to 0.75 at 5 dB, rising further to 0.87 at 0 dB, and peaking at 0.95 at -5 dB. Similarly, NSER sees an 8.4 decrease in F1 as noise levels increase from 5 to -5 dB.
| Modality | Model | MELD | IEMOCAP |
| UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | ||||||
BERT | 40.79 | 42.31 | 67.43 | 67.26 | 62.91 | 63.03 | 56.87 | 56.51 | ||||
BERT | 43.18 | 45.50 | 70.09 | 69.59 | 70.09 | 69.59 | 70.09 | 69.59 | ||||
| Speech(Whisper) | Encoder | 46.36 | 46.46 | 75.02 | 74.10 | 70.56 | 69.43 | 67.04 | 66.36 | ||||
| Speech(Whisper) | Decoder | 46.37 | 47.55 | 72.85 | 72.39 | 71.47 | 70.65 | 66.99 | 66.61 | ||||
| Speech(Whisper) | Encoder+Decoder | 48.95 | 48.87 | 76.50 | 76.53 | 73.00 | 72.42 | 68.79 | 68.39 | ||||
@@
In summary, these results emphasize the significant impact of noise on ASR performance, particularly in scenarios with complex human noise. A consistent correlation between ASR performance and NSER is evident within the same type of datasets, i.e., IEMOCAP-Human subset and IEMOCAP-Environment subset, compared with different ASR systems. Furthermore, NSER strongly depends on the type of dataset, i.e., the difference between IEMOCAPs and MELD. Moreover, as noise levels increase, both ASR models exhibit deteriorating performance on NSER tasks. This decrease is slightly less pronounced in Whisper, indicating its comparative effectiveness in environmental and human noise.
V-E Impact of ASR Representations in Speech and Text Modalities on NSER

### v4

@@ V-B Effect of Different ASR Representations
| Level | Method | MELD | IEMOCAP | IEMOCAP (Human) | IEMOCAP (Environment) | ||||||
| Park | Traffic | ||||||||||
| UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | ||
| Encoder | Last | 44.56 | 45.69 | 72.71 | 72.11 | 61.96 | 61.64 | 68.05 | 67.30 | 69.86 | 69.93 |
| Best | 44.56 | 45.69 | 72.71 | 72.11 | 63.44 | 62.61 | 68.05 | 67.30 | 69.86 | 69.93 | |
| Mean | 46.36 | 46.46 | 72.91 | 72.19 | 65.38 | 65.21 | 68.54 | 67.83 | 69.61 | 69.01 | |
| Attention | 46.65 | 46.87 | 73.39 | 73.22 | 65.70 | 65.43 | 69.51 | 69.02 | 70.74 | 70.52 | |
| Scalar Mixing | 46.69 | 47.88 | 74.25 | 74.06 | 65.94 | 65.58 | 69.87 | 69.24 | 71.64 | 71.34 | |
| Weighted Sum | 46.71 | 48.01 | 75.02 | 74.10 | 67.04 | 66.36 | 70.56 | 69.43 | 72.21 | 71.87 | |
| Decoder | Last | 45.55 | 46.55 | 70.27 | 67.91 | 60.11 | 58.81 | 65.22 | 63.43 | 66.05 | 63.91 |
| Best | 46.05 | 47.35 | 71.73 | 71.40 | 62.70 | 62.02 | 67.74 | 67.47 | 68.59 | 67.54 | |
| Mean | 45.00 | 46.58 | 70.70 | 70.29 | 62.07 | 61.42 | 65.92 | 65.62 | 66.11 | 65.93 | |
| Attention | 46.21 | 46.98 | 71.03 | 70.15 | 62.91 | 62.23 | 67.27 | 66.77 | 69.91 | 69.09 | |
| Scalar Mixing | 46.52 | 47.46 | 71.74 | 72.17 | 64.33 | 64.34 | 69.23 | 69.47 | 71.42 | 70.52 | |
| Weighted Sum | 46.37 | 47.55 | 72.75 | 72.39 | 66.99 | 66.61 | 69.72 | 68.39 | 71.90 | 69.12 | |
| ALL | Proposed | 48.95 | 48.87 | 76.50 | 76.53 | 68.79 | 68.39 | 73.00 | 72.42 | 74.39 | 74.04 |
In Experiment 2, we evaluated the impact of ASR representations for NSER under 0 dB SNR across MELD, IEMOCAP, IEMOCAP-Human, and IEMOCAP-Environment (Park, Traffic). The representations were grouped into Single-layer (last, best-performing layer) and All-layer (Mean, Attention, Scalar mixing, Weighted sum), and compared with the proposed Layer Adapter method, as detailed in Table IV.
For the MELD dataset, the best Single-layer representation (Decoder, best-performing layer) achieved 46.05% in UAR and 47.35% in F1. Among Multi-layer fusion methods, the Encoder-based Weighted Sum reached 46.71% in UAR and 48.01% in F1. In comparison, the proposed Layer Adapter achieved 48.95% in UAR and 48.87% in F1, yielding improvements of 2.90% in UAR and 1.52% in F1 over the best Single-layer, and 2.24% in UAR and 0.86% in F1 over the best Multi-layer representation.
For the IEMOCAP dataset, the best Single-layer representation (Encoder, best-performing layer) achieved 72.71% in UAR and 72.11% in F1. Among Multi-layer fusion methods, the Encoder-based Weighted Sum reached 75.02% in UAR and 74.10% in F1. In comparison, the proposed Layer Adapter attained 76.50% in UAR and 76.53% in F1, corresponding to improvements of 3.79% in UAR and 4.42% in F1 over the best Single-layer, and 1.48% in UAR and 2.43% in F1 relative to the best Multi-layer representation.
For the IEMOCAP-Human subset, the best Single-layer representation (Encoder, best-performing layer) achieved 63.44% in UAR and 62.61% in F1. The best Multi-layer fusion method (Decoder, Weighted Sum) attained 66.99% in UAR and 66.61% in F1. In contrast, the proposed method delivered 68.79% in UAR and 68.39% in F1, giving improvements of 5.35% in UAR and 5.78% in F1 over the best Single-layer, and 1.80% in UAR and 1.78% in F1 over the best Multi-layer.
For the IEMOCAP-Environment (Park), the best Single-layer representation (Decoder, best-performing layer) achieved 67.74% in UAR and 67.47% in F1. The best Multi-layer fusion approach (Encoder, Weighted Sum) reached 70.56% in UAR and 69.43% in F1. The proposed Layer Adapter achieved 73.00% in UAR and 72.42% in F1, corresponding to improvements of 5.26% in UAR and 4.95% in F1 over the best Single-layer, and 2.44% in UAR and 2.99% in F1 compared with the best Multi-layer.
For the IEMOCAP-Environment (Traffic), the best Single-layer representation (Encoder, best-performing layer) obtained 69.86% in UAR and 69.93% in F1. The strongest Multi-layer fusion method (Encoder, Weighted Sum) achieved 72.21% in UAR and 71.87% in F1. The proposed Layer Adapter further improved performance to 74.39% in UAR and 74.04% in F1, providing relative gains of 4.53% in UAR and 4.11% in F1 compared with the best Single-layer, and 2.18% in UAR and 2.17% in F1 over the best Multi-layer.
Taken together, these results demonstrate that while both Single-layer and Multi-layer fusion strategies capture valuable ASR features, the Layer Adapter method, by jointly leveraging encoder and decoder representations, consistently delivers superior performance through integrating complementary information across layers and modules.
V-C Layer-wise Analysis of ASR and SSL Representations for NSER
To further investigate how different representation levels contribute to NSER, we group the layers into three categories: low (L1–L4), middle (L5–L8), and high (L9–L12). As illustrated in Fig. 3 and Fig. 4, the performance of the ASR encoder steadily improves with increasing depth, reaching its peak in the high-layer group. In contrast, the ASR decoder achieves its best performance in the middle layers before declining. For all SSL encoders (Wav2Vec 2.0, HuBERT, and WavLM), performance improves progressively across layers and generally peaks in the high-layer group, similar to the ASR encoder.
For the IEMOCAP dataset, the ASR encoder improves steadily, starting from 60.24% at L1 and reaching a peak of 72.11% at L12 in the high-layer group. In contrast, the ASR decoder achieves its best performance in the middle layer group (71.40% at L4) before declining to 67.91% at L12. SSL models rise more gradually and reach lower peaks: Wav2Vec 2.0 attains 67.19% at L12, HuBERT 66.32% at L12, and WavLM 64.22% at L12. Even the strongest SSL model lags behind the Whisper encoder and decoder by at least 4%, highlighting that ASR representations capture richer semantic cues in clean speech.
In the Human subset, the encoder performance increases steadily from 53.40% at L1 to a peak of 62.61% at L11 in the high-layer group, whereas the decoder peaks earlier at 62.20% (L4) before declining to 58.81% at L12. Among SSL models, HuBERT performs best with 60.08% at L12, followed by WavLM (59.32%, L11), while Wav2Vec 2.0 lags behind at 57.49% (L12). Although the performance gap between the best SSL model and the Whisper encoder is narrower in this subset, SSL models still underperform overall, indicating that ASR representations remain more robust to human noise.
In the Environment–Park subset, the encoder performance increases from 54.92% at L1 to a peak of 67.30% at L12 in the high-layer group, whereas the decoder peaks earlier at 67.47% (L3) before declining to 63.43% at L12. Among SSL models, HuBERT performs best with 60.72% at L12, followed by Wav2Vec 2.0 (59.84%, L11) and WavLM (59.58%, L11). Although these results approach the decoder’s performance, they remain more than 6% below the encoder, further underscoring the superior robustness of ASR-based representations under fluctuating park noise.
In the more challenging Environment–Traffic subset, the encoder performance increases from 56.48% at L1 to a peak of 69.93% at L12 in the high-layer group, whereas the decoder peaks earlier at 67.54% (L4) and then steadily declines to 63.91% at L12. Among SSL models, Wav2Vec 2.0 performs best with 62.63% at L12, followed closely by WavLM (62.43%, L11) and HuBERT (62.13%, L11). Although these results approach the decoder’s performance, they remain more than 6% below the encoder, underscoring the superior robustness of ASR-based representations under fluctuating traffic noise.
Overall, the layer-wise analysis highlights a clear division of strengths: higher Whisper encoder layers provide dominant semantic and noise-robust features, while middle decoder layers supply complementary phonetic–semantic alignment. SSL models such as Wav2Vec 2.0, HuBERT, and WavLM exhibit gradual improvements and peak at higher layers, with HuBERT and WavLM generally outperforming Wav2Vec 2.0 in noisy subsets, yet all SSLs remain consistently below the encoder and decoder.
V-D Robustness of ASR Representations on NSER
To evaluate the robustness of ASR representations for NSER, we compared the encoder (Whisper-Enc), the decoder (Whisper-Dec), and their integration (Proposed) against state-of-the-art SSL models across different SNR levels. The results are presented in Table V.
| Model | Feature | SNR -10 dB | SNR -5 dB | SNR 0 dB | SNR 5 dB | SNR 10 dB | |||||
| UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | ||
| IEMOCAP - Human | |||||||||||
| SSL | Wav2Vec 2.0 (FT) | 45.63 | 45.22 | 50.26 | 49.26 | 59.24 | 57.49 | 64.40 | 62.66 | 65.83 | 64.80 |
| HuBERT (FT) | 51.43 | 50.22 | 54.62 | 54.57 | 61.08 | 60.08 | 64.40 | 62.90 | 65.57 | 64.48 | |
| WavLM (FT) | 53.43 | 52.03 | 56.63 | 54.33 | 59.28 | 57.36 | 61.65 | 60.89 | 64.23 | 63.92 | |
| ASR | Whisper-Enc | 61.21 | 60.54 | 61.45 | 60.77 | 67.04 | 66.36 | 70.08 | 70.22 | 71.27 | 70.87 |
| Whisper-Dec | 60.20 | 59.09 | 60.42 | 60.19 | 66.99 | 66.61 | 68.38 | 68.78 | 69.97 | 69.94 | |
| Proposed | 61.92 | 60.87 | 63.06 | 61.58 | 68.79 | 68.39 | 70.81 | 70.71 | 73.35 | 73.02 | |
| IEMOCAP - Environment (Park) | |||||||||||
| SSL | Wav2Vec 2.0 (FT) | 52.73 | 52.03 | 57.35 | 56.41 | 60.68 | 59.52 | 65.02 | 64.19 | 65.15 | 64.75 |
| HuBERT (FT) | 53.10 | 51.21 | 57.71 | 56.21 | 61.41 | 60.72 | 64.53 | 63.94 | 65.56 | 64.70 | |
| WavLM (FT) | 53.76 | 52.65 | 56.17 | 54.40 | 60.97 | 58.74 | 63.24 | 62.49 | 63.83 | 62.63 | |
| ASR | Whisper-Enc | 64.55 | 63.29 | 68.36 | 67.58 | 70.56 | 69.43 | 73.14 | 72.66 | 73.64 | 73.16 |
| Whisper-Dec | 61.36 | 59.82 | 67.20 | 65.99 | 69.72 | 68.39 | 70.07 | 68.99 | 70.42 | 69.31 | |
| Proposed | 65.37 | 64.39 | 68.78 | 67.94 | 73.00 | 72.42 | 73.84 | 73.41 | 74.83 | 74.38 | |
| IEMOCAP - Environment (Traffic) | |||||||||||
| SSL | Wav2Vec 2.0 (FT) | 55.08 | 53.06 | 60.88 | 59.17 | 64.12 | 62.63 | 65.89 | 64.88 | 66.56 | 65.59 |
| HuBERT (FT) | 55.29 | 53.93 | 59.71 | 58.59 | 62.66 | 62.04 | 65.15 | 64.09 | 67.12 | 65.86 | |
| WavLM (FT) | 55.54 | 54.23 | 59.35 | 57.69 | 62.28 | 60.95 | 65.26 | 64.15 | 65.55 | 64.41 | |
| ASR | Whisper-Enc | 67.65 | 66.82 | 71.42 | 71.29 | 72.21 | 71.87 | 72.94 | 72.81 | 73.72 | 73.25 |
| Whisper-Dec | 66.34 | 65.97 | 68.12 | 68.18 | 71.90 | 69.12 | 69.54 | 69.42 | 70.23 | 69.41 | |
| Proposed | 69.11 | 67.92 | 72.79 | 72.45 | 74.39 | 74.04 | 74.86 | 74.61 | 75.30 | 74.93 | |
In the Human subset, performance degradation is most pronounced under extreme noise. At –10 dB, the proposed method achieves 61.92% UAR and 60.87% F1. As SNR increases to 0 dB, performance improves by 6.87% in UAR and 7.52% in F1. At 10 dB, gains to –10 dB reach 11.43% in UAR and 12.15% in F1, with the method achieving 73.35% UAR and 73.02% F1. For the SSL baseline (Wav2Vec 2.0), UAR and F1 rise from 45.63% and 45.22% at –10 dB to 65.83% and 64.80% at 10 dB, corresponding to improvements of 20.20% and 19.58%. Nevertheless, at 10 dB it still trails the proposed method by 7.52% in UAR and 8.22% in F1, underscoring the robustness of encoder–decoder integration under both extremely noisy and relatively clean conditions.
In the IEMOCAP-Environment (Park) subset, at –10 dB, the proposed method achieves 65.37% UAR and 64.39% F1. As the SNR increases to 0 dB, performance improves by 7.63% in UAR and 8.03% in F1. At 10 dB, the gains relative to –10 dB reach 9.46% in UAR and 9.99% in F1, with the method achieving 74.83% UAR and 74.38% F1. For the SSL baseline (Wav2Vec 2.0), UAR and F1 increase from 52.73% and 52.03% at –10 dB to 65.15% and 64.75% at 10 dB, corresponding to improvements of 12.42% in UAR and 12.72% in F1. However, at 10 dB it still trails the proposed method by 9.68% in UAR and 9.63% in F1, confirming the superiority of the proposed approach across all noise levels.
In the IEMOCAP-Environment (Traffic) subset, at –10 dB, the proposed method achieves 69.11% UAR and 67.92% F1. As the SNR increases to 0 dB, performance improves by 5.28% in UAR and 6.12% in F1. At 10 dB, the gains relative to –10 dB reach 6.19% in UAR and 7.01% in F1, with the method achieving 75.30% UAR and 74.93% F1. For the SSL baseline (HuBERT), UAR and F1 rise from 55.29% and 53.93% at –10 dB to 67.12% and 65.86% at 10 dB, corresponding to improvements of 11.83% in UAR and 11.93% in F1. Nevertheless, at 10 dB it still falls short of the proposed method by 8.18% in UAR and 9.07% in F1, highlighting the resilience of ASR-based representations under challenging traffic noise conditions.
In addition, Fig. 5 provides a heatmap visualization of the confusion matrices across different SNR conditions, offering deeper insights into category-level robustness. A consistent observation is that Angry remains the most resilient class, with accuracy above 70% even under extreme –10 dB noise, suggesting that its prosodic and acoustic cues are less affected by background interference. In contrast, Happy exhibits the highest sensitivity to noise, often being misclassified as Neutral, particularly at lower SNR levels (–10 dB and –5 dB). This indicates that subtle emotional expressions are more vulnerable to degradation in noisy environments. Meanwhile, Sad shows relatively stable performance.
Overall, across both Human and Environmental subsets, the proposed model consistently demonstrates superior robustness compared with SSL-based features, even under extreme –10 dB noise conditions. These results confirm the proposed method as a more reliable and noise-resilient representation for NSER across diverse realistic acoustic environments.
V-E Correlation between ASR Performance and NSER Performance
@@
For the IEMOCAP dataset Whisper attains a WER of 0.28 and an NSER of 76.53 for F1. Conversely, Conformer records a WER of 0.55 and an NSER of 73.04 for F1, showing slightly lower robustness compared to Whisper.
@@
In the IEMOCAP-Human subset, performance degradation is more pronounced. For the Whisper model, at 5 dB SNR, the WER is 0.36. This WER increases to 0.57 at SNR of 0 dB and further rises to 0.86 at -5 dB, indicating a substantial decline in performance with decreasing SNR. Concurrently, the F1 score for NSER decreases by 9.13 as the noise level increases from 5 to -5 dB. For the Conformer model, performance worsens consistently across all noise levels, with the WER increasing to 0.65 at 5 dB, rising further to 0.75 at 0 dB, and peaking at 0.86 at -5 dB. Similarly, NSER sees a 6.01 decrease in F1 as noise levels increase from 5 to -5 dB.
In the IEMOCAP-Environment (Park) subset, the ASR performance gradually deteriorates as the SNR decreases. For Whisper, at an SNR of 5 dB, the WER increases to 0.32. At an SNR of 0 dB, this metric rises to 0.39. Under the most challenging condition of -5 dB, the WER peaks at 0.54. Additionally, the F1 score for NSER tasks decreases by 5.47 as the noise level increases from 5 to -5 dB, indicating degraded performance in NSER tasks under higher noise conditions. For Conformer, performance shows greater sensitivity to noise, with WER increasing to 0.75 at 5 dB SNR, rising further to 0.86 at 0 dB SNR, and reaching 0.95 at -5 dB SNR. The F1 score for NSER using the Conformer model decreases by 8.4 as the noise level increases from 5 to -5 dB, indicating that the Conformer model’s performance deteriorates under higher-noise-level conditions.
| Modality | Model | MELD | IEMOCAP | IEMOCAP (Human) | IEMOCAP (Environment) | ||||||
| Park | Traffic | ||||||||||
| UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | UAR(%) | F1(%) | ||
| Text | BERT (Whisper) | 40.79 | 42.31 | 67.43 | 67.26 | 56.87 | 56.51 | 62.91 | 63.03 | 64.65 | 64.86 |
| BERT (Transcripts) | 43.18 | 45.50 | 70.09 | 69.59 | 70.09 | 69.59 | 70.09 | 69.59 | 70.09 | 69.59 | |
| Speech | Whisper-Enc | 46.71 | 48.01 | 75.02 | 74.10 | 67.04 | 66.36 | 70.56 | 69.43 | 72.21 | 71.87 |
| Whisper-Dec | 46.37 | 47.55 | 72.75 | 72.39 | 66.99 | 66.61 | 69.72 | 68.39 | 71.90 | 69.21 | |
| Proposed | 48.95 | 48.87 | 76.50 | 76.53 | 68.79 | 68.39 | 73.00 | 72.42 | 74.39 | 74.04 | |
In the IEMOCAP-Environment (Traffic) subset, ASR performance degrades consistently with decreasing SNR. For the Whisper model, the WER increases from 0.29 at 5 dB to 0.30 at 0 dB and further to 0.41 at -5 dB. Correspondingly, the NSER F1 score decreases by 2.16 as the SNR drops from 5 to -5 dB, indicating reduced robustness under severe traffic noise. For the Conformer model, degradation is more pronounced, with the WER rising from 0.40 at 5 dB to 0.48 at 0 dB and 0.52 at -5 dB. Similarly, the NSER F1 score drops by 2.04, revealing a higher vulnerability to traffic noise compared with Whisper.
@@
In summary, these results emphasize the significant impact of noise on ASR performance, particularly in scenarios with complex human noise. A consistent correlation between ASR performance and NSER is evident within the same datasets, i.e., across the IEMOCAP-Human subset and the IEMOCAP-Environment subsets (Park and Traffic), when comparing different ASR systems. Furthermore, NSER strongly depends on the noise type, with human noise exerting a more disruptive effect than environmental noise. Moreover, as noise levels increase, both ASR models exhibit deteriorating performance on NSER tasks. This decrease is slightly less pronounced in Whisper, indicating its comparative robustness in handling diverse noisy conditions, including both human and environmental noise. Overall, aggregating across all datasets and both ASR models, a strong negative correlation between ASR performance (WER) and NSER performance (F1) is observed (Pearson’s , ).
V-F Impact of ASR Representations in Speech and Text Modalities on NSER